OpenStudy (anonymous):

help:what is the reciprocal of the number 2/7? A.-3 1/2 B. -2/7 C. 5/7 D. 3 1/2

OpenStudy (anonymous):

7/2 The reciprocal is the opposite.

hartnn (hartnn):

to get the reciprocal of a fraction, just interchange the numerator and denominator so, reciprocal of a/b is b/a now can you tell the reciprocal of 2/7 ??

OpenStudy (anonymous):

I KNOW THE RECIPOCAL IS THE OPPOSITE BUT THE CHOICES DONT MAKE SENSE TO ME :(

hartnn (hartnn):

7/2= (6+1)/2 = 6/2 +1/2 = ... ?